OpenStudy (anonymous):

Explain how to graph x + 3y = 3 using the x- and y-intercepts method. Be sure to use complete sentences and use the x- and y-intercepts method only. for 8 points

OpenStudy (anonymous):

@mathslover

OpenStudy (ivettef365):

to find the x intercept substitute the y for 0 and find value of x same for y intercept substitute the x for 0 and find value of y

OpenStudy (anonymous):

??

OpenStudy (ivettef365):

when you substitute the x for 0 then you have 0 + 3y = 3 3y = 3 y = 1 so the y-int = (0,1) now try to find the x int

OpenStudy (anonymous):

what you can also do is substitute x = 0, to give 3y = 3, y = 1, to get the first coordinate (0,1), then substitute y = 0, so x = 3, and you have another co ordinate (3,0). then join a line from one point to the other, and continue the line from both sides

OpenStudy (anonymous):

x+3(0)=3 x+0=3 x=3 correct ?

OpenStudy (anonymous):

so 3,0

OpenStudy (ivettef365):

good job, so the x-int = (3,0) now you have two points (0,1) and (3,0) to graph join the lines

OpenStudy (anonymous):

oh ok so that it ???

OpenStudy (anonymous):

if yes thank you

OpenStudy (ivettef365):

yw :)

OpenStudy (anonymous):

could you help with this as well please

OpenStudy (ivettef365):

sure, what do you need help with

OpenStudy (ivettef365):

ok, first you need to find the slope, we use formula y2-y1 ----- x2-x1

OpenStudy (ivettef365):

you have points (0,4) and (-2,8) do you know how to plug in formula

OpenStudy (ivettef365):

x1,y1 x2,y2 (0 , 4) (-2,8)